OpenStudy (anonymous):

help proving: csc^4x-cot^4x = csc^2x+cot^2x ? how to prove it?

OpenStudy (anonymous):

Write it out in terms of cosine and sines

OpenStudy (anonymous):

\[\frac{ 1 }{ \sin ^{4}x }-\frac{ \cos ^{4}x }{ \sin ^{4}x }\]

OpenStudy (anonymous):

is the lhs, and now the rhs is?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sin ^{2}x }+\frac{ \cos^{2}x }{ \sin^{2}x }\]

OpenStudy (anonymous):

so set them equal to each other and clear out the denominators by multipling through by the corresponding trigonometric functions

OpenStudy (anonymous):

(solving on paper)

OpenStudy (anonymous):

\[\frac{ \sin^{4}x-\sin^{4}xcos^{4}x }{ \sin^{4}x } = \frac{ \sin^{2}x+\sin^{2}xcos^{2}x }{ \sin^{2}x }\] ???

OpenStudy (anonymous):

Note that \[1+\cot^2x=\csc^2x\]and that:\[\csc^4x-\cot^4x=(\csc^2x-\cot^2x)(\csc^2x+\cot^2x)\]

OpenStudy (anonymous):

yes I know those identities but how will I apply it? help.

OpenStudy (anonymous):

If \(1+\cot^2x=\csc^2x\), then what is \(\csc^2x-\cot^2x\)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

nvm, go with joe's idea much neater lol

OpenStudy (anonymous):

Your done now @ECEstudent9405

OpenStudy (anonymous):

So:\[\csc^4x-\cot^4x=(\csc^2x-\cot^2x)(\csc^2x+\cot^2x)=\cdots\]

OpenStudy (anonymous):

ok. I get it sir. :)

OpenStudy (anonymous):

thanks. so it can be solved just like that. (_ _)"